# Implicit neural representations for accurate estimation of the standard model of white matter

### v1

@@ Abstract
Diffusion magnetic resonance imaging (dMRI) enables non-invasive investigation of tissue microstructure. The Standard Model (SM) of white matter aims to disentangle dMRI signal contributions from intra- and extra-axonal water compartments. However, due to the model’s high-dimensional nature, extensive acquisition protocols with multiple b-values and diffusion tensor shapes are typically required to mitigate parameter degeneracies. Even then, accurate estimation remains challenging due to noise. This work introduces a novel estimation framework based on implicit neural representations (INRs), which incorporate spatial regularization through the sinusoidal encoding of the input coordinates. The INR method is evaluated on both synthetic and in vivo datasets and compared to parameter estimates using cubic polynomials, supervised neural networks, and nonlinear least squares. Results demonstrate superior accuracy of the INR method in estimating SM parameters, particularly in low signal-to-noise conditions. Additionally, spatial upsampling of the INR can represent the underlying dataset anatomically plausibly in a continuous way, which is unattainable with linear or cubic interpolation. The INR is fully unsupervised, eliminating the need for labeled training data. It achieves fast inference ($\sim$6 minutes), is robust to both Gaussian and Rician noise, supports joint estimation of SM kernel parameters and the fiber orientation distribution function with spherical harmonics orders up to at least 8 and non-negativity constraints, and accommodates spatially varying acquisition protocols caused by magnetic gradient non-uniformities. The combination of these properties along with the possibility to easily adapt the framework to other dMRI models, positions INRs as a potentially important tool for analyzing and interpreting diffusion MRI data.
@@
Diffusion MRI , Microstructure , Deep learning , Parameter estimation , Biophysical modeling , Quantitative MRI
journal: Medical Image Analysis
@@
{highlights}
Accurate and fast parameter estimation of the Standard model of white matter
Create noise-robust continuous representations of diffusion MRI datasets
No dependency on large training datasets due to the unsupervised neural network structure
Easy incorporation of spatially varying acquisition protocols due to gradient non-uniformities
@@ 1 Introduction
Diffusion magnetic resonance imaging (dMRI) is a non-invasive technique for in-vivo measurement of water diffusion in tissue using magnetic field gradients. To extract biologically interpretable information, a common approach is to fit a microstructural tissue model to a set of signals acquired with different dMRI acquisition settings (Alexander et al., 2019; Lampinen et al., 2023; Jelescu et al., 2020). In the absence of diffusion time dependence, these typically include different combinations of gradient strengths (commonly quantified by the b-value), directions (b-vector), and B-tensor shape (Westin et al., 2014). Microstructural parameters estimated by these models – including compartmental signal fractions and diffusivities – have shown to be sensitive to changes in brain structure due to diseases like multiple sclerosis (Alotaibi et al., 2021), Alzheimer disease (Parker et al., 2018) and Parkinson’s disease (Kim et al., 2016), and can provide a more fundamental understanding of tissue microstructure in both healthy and pathological tissues (Zhang et al., 2012).
The Standard Model of white matter (SM) (Novikov et al., 2019) describes the signal arising from white matter by a kernel consisting of three compartments (intra-axonal, extra-axonal, and free water (occasionally omitted)) convolved with a fiber orientation distribution (FOD) (Tournier et al., a). Compartmental signal fractions and diffusivities can be estimated, alongside the parameters that describe the FOD (usually in the form of a spherical harmonics (SH) series). Nevertheless, the high-dimensional parameter space of the SM complicates the estimation of its parameters, potentially leading to low accuracy, precision, and degeneracy of estimates (Jelescu et al., 2016). These issues become even more prominent at high noise levels.
Multiple strategies have been employed to fit microstructure models to dMRI data. When the primary goal is to estimate the tissue’s directional structure, a common two-step approach involves first fixing the kernel using a global estimate, followed by solving a linear inverse problem to estimate the fiber orientation distribution (FOD) (Tournier et al., 2007; Jeurissen et al., 2014). In contrast, when the focus is on estimating the kernel parameters, the orientational dependence is factored out by using rotational invariants of the signal (Novikov et al., 2018; Coelho et al., 2022; Reisert et al., 2017; Kaden et al., 2016). This last approach is most common for SM parameter estimation.
@@
In this work, we implement INRs for estimation of the SM and demonstrate the noise-robustness, continuous representation, and applicability of INRs for fitting on both synthetically generated and in vivo dMRI data. Synthetically generated datasets facilitate a quantitative analysis of the model outputs, while in vivo data quantitatively shows the performance in a realistic setting. The INR is compared to two existing machine learning methods (Coelho et al., 2022; de Almeida Martins et al., 2021) for fitting the SM, as well as nonlinear least squares. Additionally, moving beyond existing methods, the FOD SH coefficients up to order eight are estimated directly alongside the SM kernel parameters. Thus, in contrast to earlier methods that focused either on the kernel or on the FOD, this approach performs a joint estimation, which can improve accuracy and ameliorate degeneracy (Henriques et al., 2025). Moreover, because this is an unsupervised method, it does not rely on large training datasets as supervised methods do. It also does not require uniform acquisition parameters across voxels, making it capable of explicitly correcting for gradient non-uniformities (Bammer et al., 2003). The latter would become impractical for supervised methods which would need prohibitively large sets of training data to capture voxel-wise protocol deviations. Altogether, the proposed method provides a flexible, noise-robust, spatially coherent way of fitting the SM, which is not biased by training data.
@@ 2.1 Standard model of white matter
Multiple approaches have been suggested to model white matter dMRI signal as a combination of sticks and anisotropic Gaussian diffusion compartments (Zhang et al., 2012; Reisert et al., 2017; Jespersen et al., 2007; Kroenke et al., 2004; Jelescu et al., 2016; Veraart et al., 2018; Fieremans et al., 2011). Generalization of this principle without introducing constraints on model parameters has led to a unified framework called the standard model of white matter (Novikov et al., 2018, 2019). The standard model assumes the measured signal $S$ to be described by the convolution of a kernel $\mathcal{K}(b,b_{\Delta},\bm{n}\cdot\bm{u})$ – describing the signal arising from water diffusing within and around a coherent fiber bundle with direction $\bm{n}$ – with a distribution of fiber populations $\mathcal{P}(\bm{n})$
$S(b,b_{\Delta},\bm{u})=S_{0}[\mathcal{K}(b,b_{\Delta},\bm{n}\cdot\bm{u})*\mathcal{P}(\bm{n})]$ (1)
@@
$\begin{split}\mathcal{K}(b,b_{\Delta},\bm{n}\cdot\bm{u})&=\exp\left(\frac{1}{3}bb_{\Delta}(D_{\parallel}-D_{\perp})-\frac{1}{3}b\left(D_{\parallel}+2D_{\perp}\right.\right.\\ &\quad\left.\left.-bb_{\Delta}\left(\bm{n}\cdot\bm{u}\right)^{2}(D_{\parallel}-D_{\perp})\right)\right)\end{split}$ (2)
@@
$S(b,b_{\Delta},\bm{u})$ $=S_{0}\cdot\Bigg{[}f_{i}\int_{|\bm{n}|=1}\exp\left(-b_{\Delta}\left(\bm{n}\cdot\bm{u}\right)^{2}D_{i}\right)$
×$\quad\times\mathcal{P}(\bm{n})\,d\bm{n}\cdot\exp\left(\frac{1}{3}bb_{\Delta}D_{i}-\frac{1}{3}bD_{i}\right)$ (3)
$+(1-f_{i})\int_{|\bm{n}|=1}\exp\left(-bb_{\Delta}\left(\bm{n}\cdot\bm{u}\right)^{2}(D_{e}-D_{p})\right)$
×$\quad\times\mathcal{P}(\bm{n})\,d\bm{n}\cdot\exp\left(\frac{1}{3}bb_{\Delta}(D_{e}-D_{p})-\frac{1}{3}b(D_{e}+2D_{p})\right)\Bigg{]}.$
@@ 2.2 INR network architecture
The purpose of the INR is to map a coordinate vector $\bm{x}$ to a desired output vector $\bm{k}$ which represents the underlying dataset at that coordinate, by passing the coordinate through a neural network $\mathcal{F}_{\Psi}:\bm{x}\rightarrow\bm{k}$ with weights $\Psi$. In our case we map a 3D-coordinate $\bm{x}\in\mathbb{R}^{3}$ to a vector of parameters for the SM kernel and FOD, $\bm{k}=[D_{i},D_{e},D_{p},f_{i},S_{0},p^{0}_{0},...p^{m}_{l}]$ where $p^{m}_{l}$ is the coefficient of the SH basis function of order $l$ and phase $m$. The signal is hence projected onto real SH as in (Lampinen et al., 2020). The implicit neural representation consists of three parts: the spatial encoding, a small multi-layer perceptron (MLP), and a number of output layers (Figure 1). Each of the parts will be discussed in-depth in the upcoming sections.
Figure 1: INR network architecture. a) shows how the input (signal + coordinates) is mapped to a higher dimensional frequency space. b) These values are then forwarded to the MLP. c) Output layer of the MLP is converted to SM parameters.
@@ 2.2.1 Spatial encoding
By encoding the input coordinates to a high dimensional space before entering them into the model, we can greatly increase the representational power of the INR (Tancik et al., 2020). We use the Fourier features encoding described by Tancik et al. (Tancik et al., 2020), which was used previously to model MSMT-CSD using INRs (Hendriks et al., 2025). First we scale the coordinates $\bm{x}$ (maintaining aspect ratio) to lie in range $[-1,1]^{3}$, and then apply to following transformation
@@ 2.2.3 Output layers
The final part of the INR architecture maps $\bm{z}$ to a parameter estimate $\bm{\hat{k}}$ using a separate fully connected layer, called ’head’, for each parameter estimate. For the SM parameters $\hat{D}_{i}$, $\hat{D}_{e}$, $\hat{D}_{p}$, and $\hat{f}_{i}$ the heads use a Sigmoid activation function scaled to fit physiological ranges (Table 1). For $\hat{S}_{0}$ a softplus activation function which ensures positivity, without an upper bound (Dugas et al., 2000). The SH-coefficients of the estimated FOD $\mathcal{\hat{P}}(\bm{n})$ require both positive and negative outputs and, therefore, have no activation function. This results in the full output layer of the INR providing the mapping $\mathcal{Q}_{\Psi_{q}}:\bm{z}\rightarrow\bm{\hat{k}}$ where $\mathcal{Q}_{\Psi_{q}}$ are the output layers with weights $\Psi_{q}$.
$D_{i}$ $D_{e}$ $D_{p}$ $f_{i}$ $S_{0}$
min. 0 0 0 0 0
max. 4 4 1.5 1 -
@@ 2.3 Model fitting
Given set of $N_{m}$ (capital $N$ denoting a fixed number, opposed to lower case hyperparameters $n_{p}$ and $n_{h}$) measurements $\{(b_{i},(b_{\Delta})_{i},\bm{u}_{i})|i\in 1,...,N_{m}\}$, measured at coordinates $\bm{x_{j}}\in\bm{X}$ with $\bm{X}\subset\mathbb{R}^{3}$ being the set of all measured coordinates in the dMRI dataset, the estimated signal $\hat{S}(b_{i},(b_{\Delta})_{i},\bm{u}_{i},\bm{x_{j}})$ at coordinate $\bm{x_{j}}$ is obtained from the model output $\mathcal{F}_{\Psi}:\bm{x}\rightarrow\bm{\hat{k}}$ by calculating the estimated kernel $\mathcal{\hat{K}}(b,b_{\Delta},\bm{n}\cdot\bm{u})$ using (2) and convolving with $\mathcal{\hat{P}}(\bm{n})$ as in (1). We approximate the desired INR $\mathcal{F}_{\Psi}$ with weights $\Psi:=\{\Psi_{m},\Psi_{q}\}$ by finding the weights $\Psi^{*}$ that minimize the error (mean squared error (MSE) or Rician likelihood loss (Parker et al., 2023), the latter allowing an explicit correction of Rician noise) between the estimated signal $\hat{S}$ and the measured signal $S$
@@
where $\mathcal{L}$ is either the MSE or the Rician likelihood loss. We include an additional term $\Lambda_{\bm{x_{j}}}$ which is a non-negativity constraint for the FOD at $\bm{x_{j}}$, as described by Tournier et al. (Tournier et al., a). The constraint is calculated by sampling the FOD across the spherical domain and adding any negative values as a loss. The fitting process is shown in Figure 2.
@@ 2.4 Implementation
The INR is implemented in Python 3.10.10 using PyTorch 2.0.0. An Adam optimizer was used with a learning rate of $10^{-4}$ and default settings otherwise. The hyperparameters were set at $n_{p}=5000$, $n_{h}=2048$, $\sigma^{2}=3.5$ for the SNR 50 synthetic datasets (see section 2.6) and the in-vivo dataset (see section 2.7), and $\sigma^{2}=2.5$ for the SNR 20 synthetic datasets. More details about the choice of $\sigma^{2}$ is given in A. Each INR was fit for 150 epochs on an NVIDIA RTX 4080 GPU with 16GB of VRAM, with a batch size of 500. Visualizations of the model output were created using matplotlib 3.8.0, and MRtrix3 3.0.4 (Tournier et al., b). Numerical integration to calculate the integral during training is implemented with Torchquad Simpson function (Gómez et al., 2021).

### v2

@@ Abstract
Diffusion magnetic resonance imaging (dMRI) enables non-invasive investigation of tissue microstructure. The Standard Model (SM) of white matter aims to disentangle dMRI signal contributions from intra- and extra-axonal water compartments. However, due to the model’s high-dimensional nature, accurately estimating its parameters poses a complex problem and remains an active field of research, in which different (machine learning) strategies have been proposed. This work introduces an estimation framework based on implicit neural representations (INRs), which incorporate spatial regularization through the sinusoidal encoding of the input coordinates. The INR method is evaluated on both synthetic and in vivo datasets and compared to existing methods. Results demonstrate superior accuracy of the INR method in estimating SM parameters, particularly in low signal-to-noise conditions. Additionally, spatial upsampling of the INR can represent the underlying dataset anatomically plausibly in a continuous way. The INR is self-supervised, eliminating the need for labeled training data. It achieves fast inference, is robust to noise, supports joint estimation of SM kernel parameters and the fiber orientation distribution function with spherical harmonics orders up to at least 8, and accommodates gradient non-uniformity corrections. The combination of these properties positions INRs as a potentially important tool for analyzing and interpreting diffusion MRI data.
@@
Diffusion MRI , Microstructure , Biophysical modeling , Quantitative MRI
@@ 1 Introduction
Diffusion magnetic resonance imaging (dMRI) is a non-invasive technique for in-vivo measurement of water diffusion in tissue using magnetic field gradients. To extract biologically interpretable information, a common approach is to fit a microstructural tissue model to a set of signals acquired with different dMRI acquisition settings (Alexander et al., 2019; Lampinen et al., 2023; Jelescu et al., 2020). In the absence of diffusion time dependence, these typically include different combinations of gradient strengths (commonly quantified by the b-value), directions (b-vector), and B-tensor shape (Westin et al., 2014). Microstructural parameters estimated by these models – including compartmental signal fractions and diffusivities – have shown to be sensitive to changes in brain structure due to diseases like multiple sclerosis (Alotaibi et al., 2021), Alzheimer’s disease (Parker et al., 2018) and Parkinson’s disease (Kim et al., 2016), and can provide a more fundamental understanding of tissue microstructure in both healthy and pathological tissues (Zhang et al., 2012).
The Standard Model of white matter (SM) (Novikov et al., 2019) describes the signal arising from white matter by a kernel consisting of three compartments (intra-axonal, extra-axonal, and free water (occasionally omitted)) convolved with a fiber orientation distribution (FOD) (Tournier et al., 2007b). Compartmental signal fractions and diffusivities can be estimated, alongside the parameters that describe the FOD (usually in the form of a spherical harmonics (SH) series). Nevertheless, the high-dimensional parameter space of the SM complicates the estimation of its parameters, potentially leading to low accuracy, precision, and degeneracy of estimates (Jelescu et al., 2016). These issues become even more prominent at high noise levels.
Multiple strategies have been employed to fit microstructure models to dMRI data. When the primary goal is to estimate the tissue’s directional structure, a common two-step approach involves first fixing the kernel using a global estimate, followed by solving a linear inverse problem to estimate the fiber orientation distribution (FOD) (Tournier et al., 2007a; Jeurissen et al., 2014). In contrast, when the focus is on estimating the kernel parameters, the orientational dependence is factored out by using rotational invariants of the signal (Novikov et al., 2018; Coelho et al., 2022; Reisert et al., 2017; Kaden et al., 2016). This last approach is most common for SM parameter estimation.
@@
Building upon our previous work (Hendriks et al., 2023, 2025), we implement INRs to estimate the SM parameters alongside the FODs, and demonstrate the noise-robustness, continuous representation, and applicability of INRs for fitting on both synthetically generated and in vivo dMRI data. Synthetically generated datasets facilitate a quantitative analysis of the model outputs, while in vivo data quantitatively shows the performance in a realistic setting. The INR is compared to two existing machine learning methods (Coelho et al., 2022; de Almeida Martins et al., 2021) for fitting the SM, as well as nonlinear least squares. Additionally, moving beyond existing methods, the FOD SH coefficients up to order eight are estimated directly alongside the SM kernel parameters. Thus, in contrast to earlier methods that focused either on the kernel or on the FOD, this approach performs a joint estimation, which can improve accuracy and ameliorate degeneracy (Henriques et al., 2025). Moreover, every INR is fit on and represents a single dMRI dataset and, therefore, does not rely on large training datasets as supervised methods do. Furthermore, it is capable of explicitly correcting for gradient non-uniformities by inputting the effective acquisition protocol (B-tensor) per voxel with the spatially varying gradient coil tensor of the scanner (Bammer et al., 2003). The latter would become impractical for supervised methods which would need prohibitively large sets of training data to capture voxel-wise protocol deviations. Altogether, the proposed method provides a flexible, noise-robust, spatially coherent way of fitting the SM, which is not biased by training data.
@@ 2.1 Standard Model of white matter
Multiple approaches have been suggested to model white matter dMRI signal as a combination of sticks and anisotropic Gaussian diffusion compartments (Zhang et al., 2012; Reisert et al., 2017; Jespersen et al., 2007; Kroenke et al., 2004; Jelescu et al., 2016; Veraart et al., 2018; Fieremans et al., 2011). Generalization of this principle without introducing constraints on model parameters has led to a unified framework called the Standard Model of white matter (Novikov et al., 2018, 2019). The Standard Model assumes the measured signal $S$ to be described by the convolution of a kernel $\mathcal{K}(b,b_{\Delta},\bm{n}\cdot\bm{u})$ – describing the signal arising from water diffusing within and around a coherent fiber bundle with direction $\bm{n}$ – with a distribution of fiber populations $\mathcal{P}(\bm{n})$ on the unit sphere
$S(b,b_{\Delta},\bm{u})=S_{0}\int_{S^{2}}\mathcal{K}(b,b_{\Delta},\bm{n}\cdot\bm{u})\,\mathcal{P}(\bm{n})\,d\bm{n}.$ (1)
@@
$\mathcal{K}_{zep}(b,b_{\Delta},\bm{n}\cdot\bm{u})=\exp\Bigg[\frac{1}{3}bb_{\Delta}(D_{\parallel}-D_{\perp})-\frac{1}{3}b\Big(D_{\parallel}+2D_{\perp}-bb_{\Delta}(\bm{n}\cdot\bm{u})^{2}(D_{\parallel}-D_{\perp})\Big)\Bigg]$ (2)
@@
$S(b,b_{\Delta},\bm{u})$ $=S_{0}\cdot\Bigg[f_{i}\int_{S^{2}}\exp\Big(-bb_{\Delta}\,(\bm{n}\cdot\bm{u})^{2}D_{i}\Big)\mathcal{P}(\bm{n})\,d\bm{n}\cdot\exp\Big(\frac{1}{3}bD_{i}-\frac{1}{3}bD_{i}\Big)$ (3)
$\quad+(1-f_{i})\int_{S^{2}}\exp\Big(-bb_{\Delta}(\bm{n}\cdot\bm{u})^{2}(D_{e}-D_{p})\Big)\mathcal{P}(\bm{n})\,d\bm{n}\cdot\exp\Big(\frac{1}{3}b(D_{e}-D_{p})-\frac{1}{3}b(D_{e}+2D_{p})\Big)\Bigg].$
@@ 2.2 INR network architecture
The purpose of the INR is to map a coordinate vector $\bm{x}$ to a desired output vector $\bm{k}$ which represents the underlying dataset at that coordinate, by passing the coordinate through a neural network $\mathcal{F}_{\Psi}:\bm{x}\rightarrow\bm{k}$ with weights $\Psi$. In our case we map a 3D-coordinate $\bm{x}\in\mathbb{R}^{3}$ to a vector of parameters for the SM kernel and FOD, $\bm{k}=[D_{i},D_{e},D_{p},f_{i},S_{0},p^{0}_{0},...p^{m}_{l}]$ where $p^{m}_{l}$ is the coefficient of the SH basis function of order $l$ and phase $m$. The signal is hence projected onto real SH as in (Lampinen et al., 2020). The end result is a representation of a single (dMRI) dataset by a neural network, from which the parameter maps can be inferred at any $\bm{x}$. The implicit neural representation consists of three parts: the spatial encoding, a small multi-layer perceptron (MLP), and a number of output layers (Figure 1). Each of the parts will be discussed in-depth in the upcoming sections.
Figure 1: INR network architecture. a) shows how the input coordinates are mapped to a higher dimensional frequency space. b) These values are then forwarded to the MLP. c) Output layer of the MLP is converted to SM parameters.
@@ 2.2.1 Spatial encoding
By encoding the input coordinates to a high dimensional space before entering them into the model, we can greatly increase the representational power of the INR (Tancik et al., 2020). We use the Fourier features encoding described by Tancik et al. (Tancik et al., 2020), which was used previously to model MSMT-CSD using INRs (Hendriks et al., 2025). First we scale the coordinates $\bm{x}$ (maintaining aspect ratio) to lie in range $1,1]^{3}$, and then apply to following transformation
@@ 2.2.3 Output layers
The final part of the INR architecture maps $\bm{z}$ to a parameter estimates $\bm{\hat{k}}$ using a separate fully connected layer, called ’head’, for each parameter estimate. For the SM parameters $\hat{D}_{i}$, $\hat{D}_{e}$, $\hat{D}_{p}$, and $\hat{f}_{i}$ the heads use a Sigmoid activation function scaled to fit physiological ranges (Table 1). For $\hat{S}_{0}$ a softplus activation function was used which ensures positivity, without an upper bound (Dugas et al., 2000). The SH-coefficients of the estimated FOD $\mathcal{\hat{P}}(\bm{n})$ require both positive and negative outputs and, therefore, have no activation function. This results in the full output layer of the INR providing the mapping $\mathcal{Q}_{\Psi_{q}}:\bm{z}\rightarrow\bm{\hat{k}}$ where $\mathcal{Q}_{\Psi_{q}}$ are the output layers with weights $\Psi_{q}$.
$f_{i}$ $D_{i}$ $D_{e}$ $D_{p}$ $S_{0}$ $p_{l}^{m}$
min. 0 0 0 0 0 -inf.
max. 1 4 4 1.5 +inf. +inf.
@@ 2.3 Model fitting
Given set of $N_{m}$ (capital $N$ denoting a fixed number, opposed to lower case hyperparameters $n_{p}$ and $n_{h}$) measurements $\{(b_{i},(b_{\Delta})_{i},\bm{u}_{i})|i\in 1,...,N_{m}\}$, measured at coordinates $\bm{x_{j}}\in\bm{X}$ with $\bm{X}\subset\mathbb{R}^{3}$ being the set of all measured coordinates in the dMRI dataset, the estimated signal $\hat{S}(b_{i},(b_{\Delta})_{i},\bm{u}_{i},\bm{x_{j}})$ at coordinate $\bm{x_{j}}$ is obtained from the model output $\mathcal{F}_{\Psi}:\bm{x}\rightarrow\bm{\hat{k}}$ by calculating the estimated kernel $\mathcal{\hat{K}}(b,b_{\Delta},\bm{n}\cdot\bm{u})$ using (2) and convolving with $\mathcal{\hat{P}}(\bm{n})$ as in (1). We approximate the desired INR $\mathcal{F}_{\Psi}$ with weights $\Psi:=\{\Psi_{m},\Psi_{q}\}$ by finding the weights $\Psi^{*}$ that minimize the error (mean squared error (MSE) or Rician likelihood loss (Parker et al., 2025), the latter allowing an explicit correction of Rician noise) between the estimated signal $\hat{S}$ and the measured signal $S$
@@
where $\mathcal{L}$ is either the MSE or the Rician likelihood loss. We include an additional term $\Lambda_{\bm{x_{j}}}$ which is a non-negativity constraint for the FOD at $\bm{x_{j}}$, as described by Tournier et al. (Tournier et al., 2007b). The constraint is calculated by sampling the FOD across the spherical domain and adding any negative values as a loss. The fitting process is shown in Figure 2.
@@ 2.4 Implementation
The INR is implemented in Python 3.10.10 using PyTorch 2.0.0. An Adam optimizer was used with a learning rate of $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, and no weight decay. The hyperparameters were set at $n_{p}=5000$, $n_{h}=2048$, $\sigma^{2}=3.5$ for the SNR 50 synthetic datasets (see section 2.6) and the in-vivo dataset (see section 2.7), and $\sigma^{2}=2.5$ for the SNR 20 synthetic datasets. More details about the choice of $\sigma^{2}$ is given in the supplementary information. Each INR was fit for 150 epochs on an NVIDIA RTX 4080 GPU with 16GB of VRAM, with a batch size of 500. Visualizations of the model output were created using matplotlib 3.8.0, and MRtrix3 3.0.4 (Tournier et al., 2019). Numerical integration to calculate the integral during training is implemented with Torchquad Simpson function (Gómez et al., 2021).